# Optimal and Fair Encouragement Policy Evaluation and Learning

### v4

@@ C.2.4 Best-response oracles
$\begin{aligned} L(\pi,{\lambda})&=\hat{V}(\pi)+{\lambda}^{\top}({M}\hat{{h}}(\pi)-\hat{{d}})\\ &=\hat{V}(\pi)-{\lambda}^{\top}\hat{{d}}+\sum_{k,j}\frac{M_{k,j}\lambda_{k}}{p_{j}}{\mathbb{E}}_{n}\left[g_{j}(O,\pi){1}\left\{O\in\mathcal{E}_{j}\right\}\right].\end{aligned}$
@@
$\operatorname{BEST}_{\boldsymbol{\lambda}}(Q)$ is the best response of the $\Lambda$ player. It can be chosen to be either 0 or put all the mass on the most violated constraint. Let ${{\gamma}}(Q):={M}{{h}}(Q)$ denote the constraint values, then $\operatorname{BEST}_{\boldsymbol{\lambda}}(Q)$ returns
$\begin{cases}\mathbf{0}&\text{ if }\widehat{\gamma}(Q)\leq\widehat{\mathbf{c}}\\ B\mathbf{e}_{k^{*}}&\text{ otherwise, where }k^{*}=\arg\max_{k}\left[\widehat{\gamma}_{k}(Q)-\widehat{c}_{k}\right]\end{cases}$
C.2.5 Weighted classification reduction
@@
$L(\theta)=\mathbb{E}[\left|\psi\right|\ell(g_{\theta}(X),\operatorname{sgn}(\psi))]$
@@
Example 8 (Treatment parity, continued (weighted classification reduction)).
@@
$L(\beta)=\mathbb{E}\left[|\tilde{\psi}^{\lambda}|\ell\left(f_{\beta}(X),\operatorname{sgn}(\tilde{\psi}^{\lambda})\right)\right],\qquad\text{ where }\tilde{\psi}^{\lambda}=\psi+\frac{\lambda_{A}}{p_{A}}(p_{1\mid 1}-p_{1\mid 0})-\sum_{a\in\mathcal{A}}\lambda_{a}.$
@@
Example 9 (Responder-conditional treatment parity, continued).
The Lagrangian is $L(\beta)=\mathbb{E}\left[|\tilde{\psi}^{\lambda}|\ell\left(f_{\beta}(X),\operatorname{sgn}(\tilde{\psi}^{\lambda})\right)\right]$ with weights:
$\tilde{\psi}^{\lambda}=\psi+\frac{\lambda_{A}}{p_{A}}\frac{(p_{1\mid 1}-p_{1\mid 0})(\mu_{1}-\mu_{0})}{\mathbb{E}_{n}[(\mu_{1}(X,A)-\mu_{0}(X,A))\mid A=a]}-\sum_{a\in\mathcal{A}}\lambda_{a}.$
where $p_{a}:=\hat{P}(A=a)$ and $\lambda_{a}:=\lambda_{(a,+)}-\lambda_{(a,-)}$.
C.3 Proofs
Proof of Proposition˜2.
$\mathbb{E}[c(\pi,T(\pi),Y(\pi))]$ $=\textstyle\sum_{t\in\mathcal{T},r\in\{0,1\}}\mathbb{E}[\pi_{r}(X)\mathbb{E}[\mathbb{I}\left[T(r)=t\right]c_{rt}(Y(t(r)))\mid R=r,X]]$
$\textstyle=\sum_{t\in\mathcal{T},r\in\{0,1\}}\mathbb{E}[\pi_{r}(X){P}(T=t\mid R=r,X)\mathbb{E}[c_{rt}(Y(t(r)))\mid R=r,X]]$
$\textstyle=\sum_{t\in\mathcal{T},r\in\{0,1\}}\mathbb{E}[\pi_{r}(X){P}(T=t\mid R=r,X)\mathbb{E}[c_{rt}(Y)\mid T=t,X]]$
@@ C.3 Proofs
where the last line follows by the conditional exclusion restriction (Assumption˜2) and consistency (Assumption˜1).
Proof of Proposition˜8.
$P(T(\pi)=1\mid A=a,Y(1(\pi))>Y(0(\pi)))$
$=\frac{P(T(\pi)=1,Y(1(r))>Y(0(r))\mid A=a)}{P(Y(1(\pi))>Y(0(\pi))\mid A=a)}$ by Bayes’ rule
$=\frac{P(T(\pi)=1,Y(1)>Y(0)\mid A=a)}{P(Y(1)>Y(0)\mid A=a)}$ by Assumption˜2
$=\frac{\sum_{r}\mathbb{E}[\mathbb{E}[\pi_{r}(X)\mathbb{I}\left[T(r)=1\right]\mathbb{I}\left[Y(1)>Y(0)\right]\mid A=a,X]])}{P(Y(1)>Y(0)\mid A=a)}$ by iter. exp
$\frac{\sum_{r}\mathbb{E}[\pi_{r}(X)p_{1\mid r}(X,A)(\mu_{1}(X,A)-\mu_{0}(X,A))\mid A=a])}{\mathbb{E}[(\mu_{1}(X,A)-\mu_{0}(X,A))\mid A=a]}$ by Proposition˜2
Appendix D Proofs
D.1 Proofs for generalization under unconstrained policies
Proposition 9 (Policy value generalization).
@@ D.1 Proofs for generalization under unconstrained policies
Assume the nuisance models $\eta=[p_{1\mid 0},p_{1\mid 1},\mu_{1},\mu_{0},e_{r}(X)]^{\top},\eta\in\mathcal{F}_{\eta}$ are consistent and well-specified with finite VC-dimension $v_{\eta}$ over the product function class $H$. We provide a proof for the general case, including doubly-robust estimators, which applies to the statement of Proposition˜9 by taking $\eta=[p_{1\mid 0},p_{1\mid 1},\mu_{1},\mu_{0}].$
Let $\Pi=\{\mathbb{I}\{\mathbb{E}[L(\lambda,X,A;\eta)\mid X]<0\}\colon\lambda\in\mathbb{R};\eta\in\mathcal{F}_{\eta}\}.$
$\sup_{\pi\in\Pi,\lambda\in\mathbb{R}}\left|(\mathbb{E}_{n}[\pi L(\lambda,X,A)]-\mathbb{E}[\pi L(\lambda,X,A)])\right|=O_{p}(n^{-\frac{1}{2}})$
@@
Corollary 5.
$\mathbb{E}[L(\hat{\lambda},X,A)_{+}]\leq\mathbb{E}[L(\lambda^{*},X,A)_{+}]+O_{p}(n^{-\frac{1}{2}})$
Proof of Proposition˜9.
We study a general Lagrangian, which takes as input pseudo-outcomes $\psi^{t\mid r}(O;\eta),\psi^{y\mid t}(O;\eta),\psi^{1\mid 0,\Delta A}$ where each satisfies that
@@
We make high-level stability assumptions on pseudooutcomes $\psi$ relative to the nuisance functions $\eta$ (these are satisfied by standard estimators that we will consider):
Assumption 11.
$\psi^{t\mid r},\psi^{y\mid t},\psi^{1\mid 0,\Delta A}$ respectively are Lipschitz contractions with respect to $\eta$ and bounded
@@
$L(\lambda,X,A;\eta)=\psi_{t\mid r}(O;\eta)\left\{\psi_{y\mid t}(O;\eta)+\frac{\lambda}{p(A)}(\mathbb{I}\left[A=a\right]-\mathbb{I}\left[A=b\right])\right\}+\lambda(\psi^{1\mid 0,\Delta A}(O;\eta))$
@@
$\sup_{\pi\in\Pi,\lambda\in\mathbb{R}}\left|(\mathbb{E}_{n}[\pi L(\lambda,X,A)]-\mathbb{E}[\pi L(\lambda,X,A)])\right|=O_{p}(n^{-\frac{1}{2}})$
@@
$\mathbb{E}_{n}[\pi L(\lambda,X,A)]=\mathbb{E}[\pi L(\lambda,X,A)])+O_{p}(n^{-\frac{1}{2}})$
Write Lagrangian as
$\max_{\pi}\min_{\lambda}=\min_{\lambda}\max_{\pi}=\min_{\lambda}\mathbb{E}[L(O,\lambda;\eta)_{+}]$
@@
Applying Equation˜5 to each of the component classes $\mathcal{F}_{L_{1}}(O_{1:n}),\mathcal{F}_{L_{2}}(O_{1:n}),\mathcal{F}_{L_{3}}(O_{1:n}),$ and Lipschitz contraction [Bartlett and Mendelson, 2002, Thm. 12.4] of the positive part function $\mathcal{F}_{L_{+}}$, we obtain the bound
$\sup_{\lambda,\eta}\left|\mathbb{E}_{n}[L(O,\lambda;\eta)_{+}]-\mathbb{E}[L(O,\lambda;\eta)_{+}]\right|\leq\sqrt{2}(C^{L}_{t\mid r}C^{L}_{y\mid t}+C^{L}_{t\mid r}B_{p_{a}}B+BC^{L}_{1\mid 0,\Delta A})\sum_{k}\mathcal{R}(H_{k})$
Proposition 10 (Threshold solutions).
Define
$L(\lambda,X,A)=(p_{1\mid 1}(X,A)-p_{1\mid 0}(X,A))\left\{\tau(X,A)+\frac{\lambda}{p(A)}(\mathbb{I}\left[A=a\right]-\mathbb{I}\left[A=b\right])\right\}+\lambda(p_{1\mid 0}(X,a)-p_{1\mid 0}(X,b))$
$\lambda^{*}\in\arg\max_{\lambda}\;\mathbb{E}[L(\lambda,X,A)_{+}],\;\;\pi^{*}(x,u)=\mathbb{I}\{L(\lambda^{*},X,u)<0\}$
If instead $d(x)$ is a function of covariates $x$ only,
$\lambda^{*}\in\arg\max_{\lambda}\;\mathbb{E}[\mathbb{E}[L(\lambda,X,A)\mid X]_{+}],\;\;\pi^{*}(x)=\mathbb{I}\{\mathbb{E}[L(\lambda^{*},X,A)\mid X]<0\}$
Proof of Proposition˜10.
The characterization follows by strong duality in infinite-dimensional linear programming [Shapiro, 2001]. Strict feasibility can be satisfied by, e.g. solving eq.˜2 to set ranges for $\epsilon$. ∎
D.2 Proofs for robust characterization
Proof of Proposition˜7.
$V(\pi)=$ $\sum_{t\in\mathcal{T},r\in\{0,1\}}\mathbb{E}[\pi_{r}(X)\mathbb{E}[c_{rt}(Y(t))\mathbb{I}\left[T(r)=t\right]\mid R=r,X]]$
$=$ $\sum_{t\in\mathcal{T},r\in\{0,1\}}\mathbb{E}[\pi_{r}(X)\mathbb{E}[c_{rt}(Y(t))\mid R=r,X]P(T(r)=t\mid R=r,X)]$ unconf.
$=$ $\sum_{t\in\mathcal{T},r\in\{0,1\}}\mathbb{E}[\pi_{r}(X)\mathbb{E}[c_{rt}(Y(t))\mid X]P(T(r)=t\mid R=r,X)]$ Assumption˜2 (ER) (6)
$=$ $\sum_{t\in\mathcal{T},r\in\{0,1\}}\mathbb{E}\left[\pi_{r}(X)\mathbb{E}\left[c_{rt}(Y(t))\frac{\mathbb{I}\left[T(r)=t\right]}{p_{t}(X)}\mid X\right]P(T(r)=t\mid R=r,X)\right]$ unconf.
$=$ $\sum_{t\in\mathcal{T},r\in\{0,1\}}\mathbb{E}\left[\pi_{r}(X)\left\{\mathbb{E}\left[c_{rt}(Y(t))\frac{\mathbb{I}\left[T(r)=t\right]}{p_{t}(X)}+\left(1-\frac{T}{p_{t}(X)}\right)\mu_{t}(X)\mid X\right]p_{t\mid r}(X)\right\}\right]$ control variate
$=$ $\sum_{t\in\mathcal{T},r\in\{0,1\}}\mathbb{E}\left[\pi_{r}(X)\left\{\left\{c_{y}{rt}(Y(t))\frac{\mathbb{I}\left[T(r)=t\right]}{p_{t}(X)}+\left(1-\frac{T}{p_{t}(X)}\right)\mu_{t}(X)\right\}p_{t\mid r}(X)\right\}\right]$ (LOTE)
where $p_{t}(X)=P(T=t\mid X)$ (marginally over $R$ in the observational data) and (LOTE) is an abbreviation for the law of total expectation. ∎
Proof of Lemma˜1.

### v5

@@ 11.2.4 Best-response oracles
$\mathcal{L}(\pi,\lambda)=\hat{V}(\pi)+\lambda^{\top}(\hat{\Delta}(\pi)-d)$
@@
$\operatorname{BEST}_{\bm{\lambda}}(Q)$ is the best response of the $\Lambda$ player. It can be chosen to be either 0 or put all the mass on the most violated constraint. $\operatorname{BEST}_{\bm{\lambda}}(Q)$ returns
$\begin{cases}\mathbf{0}&\text{ if }\widehat{\Delta}(Q)\leq\widehat{\mathbf{c}}\\ B\mathbf{e}_{k^{*}}&\text{ otherwise, where }k^{*}=\arg\max_{k}\left[\widehat{\Delta}_{k}(Q)-\widehat{c}_{k}\right]\end{cases}$
11.2.5 Weighted classification reduction
@@
$\mathcal{L}(\theta)=\mathbb{E}[\left|\psi\right|\ell(g_{\theta}(X),\operatorname{sgn}(\psi))]$
@@
Example 11.2 (Treatment parity, continued (weighted classification reduction))
@@
$\mathcal{L}(\beta)=\mathbb{E}\left[|\tilde{\psi}^{\lambda}|\ell\left(f_{\beta}(X),\operatorname{sgn}(\tilde{\psi}^{\lambda})\right)\right],\qquad\text{ where }\tilde{\psi}^{\lambda}=\psi+\frac{\lambda_{A}}{p_{A}}(p_{1\mid 1}-p_{1\mid 0})-\sum_{a\in\mathcal{A}}\lambda_{a}.$
@@
12 Proofs
12.1 Identification, estimation and characterization
Proof 12.1
Proof of Proposition 3.1
$\mathbb{E}[Y(\pi)]$ $=\textstyle\sum_{t\in\mathcal{T},r\in\{0,1\}}\mathbb{E}[\pi_{r}(X)\mathbb{E}[\mathbb{I}\left[T(r)=t\right]Y(t(r))\mid R=r,X]]$
$\textstyle=\sum_{t\in\mathcal{T},r\in\{0,1\}}\mathbb{E}[\pi_{r}(X){P}(T=t\mid R=r,X)\mathbb{E}[Y(t(r))\mid R=r,X]]$
$\textstyle=\sum_{t\in\mathcal{T},r\in\{0,1\}}\mathbb{E}[\pi_{r}(X){P}(T=t\mid R=r,X)\mathbb{E}[Y\mid T=t,X]]$
@@ 12.1 Identification, estimation and characterization
where the last line follows by the conditional exclusion restriction (Section 3) and consistency (Section 3).
Proof 12.2
Proof of Proposition 9.1 We will repeatedly use a simple lemma and Gaussian probit identity. Generically, if $G\sim N(m,v)$, then
$\mathbb{E}[\Phi(G)]=\Phi\left(\frac{m}{\sqrt{1+v}}\right)$
Indeed, if $Z\sim N(0,1)$ is independent of $G$, then
$\mathbb{E}[\Phi(G)]=\mathbb{E}[P(Z\leq G)]=\mathbb{E}[P(Z\leq G\mid G)]=\operatorname{Pr}(Z\leq G)=\operatorname{Pr}(G-Z\leq 0)=\Phi\left(\frac{m}{\sqrt{1+v}}\right).$
First we consider $\mathbb{E}[\kappa(X,A)\mid X^{s},A=g]$. Note that
$\mathbb{E}[\kappa(X,A)\mid X^{s},A=g]=\theta^{\kappa}\mu_{g}+\theta^{s}X^{s},\qquad\;\;Var[\kappa(X,A)\mid X^{s},A=g]=(\theta^{\kappa})^{2}(\sigma_{x}^{\kappa})^{2}$
Analogously, for $\mathbb{E}[\kappa(X,A)\mid A=g]:$
$\mathbb{E}[\kappa(X,A)\mid A=g]=\theta^{\kappa}\mu_{g},\qquad\;\;Var[\kappa(X,A)\mid X^{s},A=g]=(\theta^{\kappa})^{2}(\sigma_{x}^{\kappa})^{2}+(\theta^{s})^{2}(\sigma_{x}^{s})^{2}$
In the model, $(\sigma_{x}^{\kappa})^{2},(\sigma_{x}^{s})^{2}=1$, yielding the result.
Bounding $V(\pi^{*})-V(\pi_{DP}^{blind})$: For the policy bound, note that
$V(\pi^{*})-V(\pi_{DP}^{blind})$ $=\mathbb{E}[(\pi^{*}(X,A)-\pi_{DP}^{blind}(X^{s},X^{\tau})s(X,A)]$
$=\mathbb{E}[\pi^{*}(X,A)s(X,A)]-\mathbb{E}[\pi_{DP}^{blind}(X^{s},X^{\tau})s(X^{s},X^{\tau})]$
$=\mathbb{E}[\pi^{*}(X,A)(s(X,A)-s(X^{s},X^{\tau}))]+\mathbb{E}[(\pi^{*}(X,A)-\pi_{DP}^{blind}(X^{s},X^{\tau}))s(X^{s},X^{\tau})]$
$\leq\mathbb{E}[\pi^{*}(X,A)(s(X,A)-s(X^{s},X^{\tau}))]$ (by suboptimality of $\pi^{*}$ w.r.t. $(X^{s},X^{\tau})$)
$\leq\mathbb{E}[\left|(s(X,A)-s(X^{s},X^{\tau}))\right|]$ (since $\pi^{*}\in[0,1]$)
Next we can apply results specific to the Gaussian model of Equation 5. By Lipschitzness of the Gaussian probit/cdf, $\Phi^{\prime}(x)\leq\frac{1}{\sqrt{2\pi}}.$ Therefore:
$\mathbb{E}[\left|s(X,A)-s(X^{s},X^{\tau})\right|]$ $\leq\frac{(1-\bar{p}_{0})\mathbb{E}[\left|\tau(X^{s},X^{\tau})(\theta^{\kappa}X^{\kappa}-\mathbb{E}[\theta^{\kappa}X^{\kappa}\mid X^{s},X^{\tau}])\right|}{\sqrt{2\pi}}$
$=\frac{(1-\bar{p}_{0})\mathbb{E}[\left|\tau(X^{s},X^{\tau})\right|]\mathbb{E}[\left|(\theta^{\kappa}X^{\kappa}-\mathbb{E}[\theta^{\kappa}X^{\kappa}\mid X^{s},X^{\tau}])\right|}{\sqrt{2\pi}}$
$\leq\frac{(1-\bar{p}_{0})\left|\theta^{\kappa}\right|\mathbb{E}[\left|\tau(X^{s},X^{\tau})\right|]\mathbb{E}[\left|(X^{\kappa}-\mathbb{E}[X^{\kappa}])\right|}{\sqrt{2\pi}}.$
12.2 Estimation - proofs for generalization under unconstrained policies
Proposition 12.3 (Policy value generalization)
@@ 12.2 Estimation - proofs for generalization under unconstrained policies
Assume the nuisance models $\eta=[p_{1\mid 0},p_{1\mid 1},\mu_{1},\mu_{0},e_{r}(X)]^{\top},\eta\in\mathcal{F}_{\eta}$ are consistent and well-specified with finite VC-dimension $v_{\eta}$ over the product function class $H$. We provide a proof for the general case, including doubly-robust estimators, which applies to the statement of Proposition 12.3 by taking $\eta=[p_{1\mid 0},p_{1\mid 1},\mu_{1},\mu_{0}].$
Let $\Pi=\{\mathbb{I}\{\mathbb{E}[\ell(\lambda,X,A;\eta)\mid X]<0\}\colon\lambda\in\mathbb{R};\eta\in\mathcal{F}_{\eta}\}.$
$\sup_{\pi\in\Pi,\lambda\in\mathbb{R}}\left|(\mathbb{E}_{n}[\pi\ell(\lambda,X,A)]-\mathbb{E}[\pi\ell(\lambda,X,A)])\right|=O_{p}(n^{-\frac{1}{2}})$
@@
Corollary 12.4
$\mathbb{E}[\ell(\hat{\lambda},X,A)_{+}]\leq\mathbb{E}[\ell(\lambda^{*},X,A)_{+}]+O_{p}(n^{-\frac{1}{2}})$
Proof 12.5
Proof of Proposition 12.3 We study a general Lagrangian, which takes as input pseudo-outcomes $\psi^{t\mid r}(O;\eta),\psi^{y\mid t}(O;\eta),\psi^{1\mid 0,\Delta A}$ where each satisfies that
@@
We make high-level stability assumptions on pseudooutcomes $\psi$ relative to the nuisance functions $\eta$ (these are satisfied by standard estimators that we will consider): {assumption} $\psi^{t\mid r},\psi^{y\mid t},\psi^{1\mid 0,\Delta A}$ respectively are Lipschitz contractions with respect to $\eta$ and bounded
@@
$\ell(\lambda,X,A;\eta)=\psi_{t\mid r}(O;\eta)\left\{\psi_{y\mid t}(O;\eta)+\frac{\lambda}{p(A)}(\mathbb{I}\left[A=a\right]-\mathbb{I}\left[A=b\right])\right\}+\lambda(\psi^{1\mid 0,\Delta A}(O;\eta))$
@@
$\sup_{\pi\in\Pi,\lambda\in\mathbb{R}}\left|(\mathbb{E}_{n}[\pi\ell(\lambda,X,A)]-\mathbb{E}[\pi\ell(\lambda,X,A)])\right|=O_{p}(n^{-\frac{1}{2}})$
@@
$\mathbb{E}_{n}[\pi\ell(\lambda,X,A)]=\mathbb{E}[\pi\ell(\lambda,X,A)])+O_{p}(n^{-\frac{1}{2}})$
Write the pointwise-score characterization as
$\max_{\pi}\min_{\lambda}=\min_{\lambda}\max_{\pi}=\min_{\lambda}\mathbb{E}[\ell(O,\lambda;\eta)_{+}]$
@@
Applying Equation 10 to each of the component classes $\mathcal{F}_{L_{1}}(O_{1:n}),\mathcal{F}_{L_{2}}(O_{1:n}),\mathcal{F}_{L_{3}}(O_{1:n}),$ and Lipschitz contraction [Bartlett and Mendelson, 2002, Thm. 12.4] of the positive part function $\mathcal{F}_{L_{+}}$, we obtain the bound
$\sup_{\lambda,\eta}\left|\mathbb{E}_{n}[\ell(O,\lambda;\eta)_{+}]-\mathbb{E}[\ell(O,\lambda;\eta)_{+}]\right|\leq\sqrt{2}(C^{L}_{t\mid r}C^{L}_{y\mid t}+C^{L}_{t\mid r}B_{p_{a}}B+BC^{L}_{1\mid 0,\Delta A})\sum_{k}\mathcal{R}(H_{k})$
Proof 12.6
Proof of Proposition 5.2 By Proposition 3.1, expanding $\sum_{t,r}\pi_{r}\mu_{t}p_{t\mid r}$ in the binary case and using $\kappa(X,A)\tau(X,A)=s(X,A)$ (the encouragement-score factoring), $\mathbb{E}[Y(\pi)]=C_{Y}+\mathbb{E}[\pi(X,A)s(X,A)]$ where $C_{Y}$ is independent of $\pi$. Similarly, $T(\pi)=\pi(X,A)\kappa(X,A)+p_{1\mid 0}(X,A)$, so $\mathbb{E}[T(\pi)\mid A=u]=\mathbb{E}[\pi(X,A)\kappa(X,A)+p_{1\mid 0}(X,A)\mid A=u]$ and
$\Delta_{T}(\pi):=\mathbb{E}[T(\pi)\mid A=a]-\mathbb{E}[T(\pi)\mid A=b]=\mathbb{E}[\pi(X,A)\kappa(X,A)g(A)]+c_{0}.$
Hence the Lagrangian for (4) is
$L(\pi,\lambda)=C_{Y}+\lambda(\epsilon-c_{0})+\mathbb{E}\!\left[\pi(X,A)\ell(\lambda,X,A)\right],\qquad\lambda\geq 0.$
For fixed $\lambda$, the integrand is linear in $\pi(x,u)\in\{0,1\}$, so the pointwise maximizer is $\pi_{\lambda}(x,u)=\mathbb{I}\{\ell(\lambda,x,u)>0\}$, giving $\sup_{\pi}L(\pi,\lambda)=C_{Y}+\lambda(\epsilon-c_{0})+\mathbb{E}[\ell(\lambda,X,A)_{+}]$. By strong duality for infinite-dimensional linear programs [Shapiro, 2001], any $\lambda^{*}$ minimizing this dual objective yields an optimal $\pi^{*}=\pi_{\lambda^{*}}$. If $\pi$ depends only on $X$, then $\mathbb{E}[\pi(X)\ell(\lambda,X,A)]=\mathbb{E}[\pi(X)\mathbb{E}[\ell(\lambda,X,A)\mid X]]$, so the same pointwise argument gives $\pi_{\lambda}(x)=\mathbb{I}\{\mathbb{E}[\ell(\lambda,X,A)\mid X=x]>0\}$ with dual objective $\mathbb{E}[\mathbb{E}[\ell(\lambda,X,A)\mid X]_{+}]+\lambda(\epsilon-c_{0})$. Strict feasibility can be verified using eq. 7.
12.3 Proofs for robust characterization
12.3.1 Additional results
Lemma 12.7 (Binary outcomes, constant bound)